OpenStudy (anonymous):

Calculate the circumference of the circles with the following dimensions. C = circumference D = diameter R = radius PI = 3.14 D = 4 i

OpenStudy (anonymous):

Calculate the circumference of the circles with the following dimensions. C = circumference D = diameter R = radius PI = 3.14 D = 4 in D = 6 in D = 12 i

OpenStudy (mysesshou):

\[C=2*\pi*R = \pi*D\] 1) C=3.14 * 4 =12.56 inches 2) C=3.14*6 = 18.84 inches 3) C=3.14*12 = 37.68 inches